OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (freckles):

what it is the common difference that is what are the following: 5-(-4)=? 14-5=? 23-14=?

OpenStudy (anonymous):

9 @freckles

OpenStudy (freckles):

right and we know a arithmetic sequence is of the form: \[a_n=a_1+d(n-1) \\ \text{ where } a_1 \text{ is first term } \\ \text{ and } d \text{ is common difference } \\ \text{ your answer will be of the form } \sum_{i=1}^{n} a_i=\sum_{i=1}^n [a_1+d(i-1)]\] so the only thing left to figure out is the n for which you get the last term 131

OpenStudy (freckles):

for what n is a_n =131 well just solve: \[131=a_1+d(n-1)\] where you found d to be 9

OpenStudy (anonymous):

wait so what do i do next? @freckles

OpenStudy (freckles):

did you find n yet?

OpenStudy (freckles):

for when an is 131

OpenStudy (freckles):

because that is basically the last step besides to plug into the final form

OpenStudy (anonymous):

No I didn't find it

OpenStudy (freckles):

ok well do you still need any help?

OpenStudy (freckles):

I'm not sure if that means you are having a problem solving the equation I gave you or not